### v1

# Disturbance Compensation for Safe Kinematic Control of Robotic Systems with Closed Architecture

Abstract
In commercial robotic systems, it is common to encounter a closed inner-loop torque controller that is not user-modifiable. However, the outer-loop controller, which sends kinematic commands such as position or velocity for the inner-loop controller to track, is typically exposed to users. In this work, we focus on the development of an easily integrated add-on at the outer-loop layer by combining disturbance rejection control and robust control barrier function for high-performance tracking and safe control of the whole dynamic system of an industrial manipulator. This is particularly beneficial when 1) the inner-loop controller is imperfect, unmodifiable, and uncertain; and 2) the dynamic model exhibits significant uncertainty. Stability analysis, formal safety guarantee proof, and hardware experiments with a PUMA robotic manipulator are presented. Our solution demonstrates superior performance in terms of simplicity of implementation, robustness, tracking precision, and safety compared to the state of the art. Video: https://youtu.be/zw1tanvrV8Q
I Introduction
Robotic systems typically use a hierarchical architecture combining perception, planning, and low-level control, which improves modularity and reliability. In the control layer, a kinematic controller (outer loop) and a dynamic controller (inner loop) are commonly used. Since tuning the inner loop requires expertise, it is usually hidden from users for safety, forming a closed architecture [1, 2, 3, 4]. Users can thus only design the kinematic controller, sending position or velocity commands. However, wear, modifications, or payload changes can alter dynamics and degrade inner-loop performance [5]. This work addresses the challenge of achieving unified high-performance and safe tracking control using only the kinematic controller, despite an imperfect, unmodifiable inner loop, and uncertain dynamics.
Fig. 1 shows that our approach successfully follows an infinity-shaped reference trajectory despite various uncertainties and a poor inner-loop unknown controller (see the blue trajectory). In contrast, using the poor inner-loop controller without our approach’s compensation results in poor trajectory tracking (see the red trajectory).
Figure 1: Trajectory tracking of the PUMA 500 for an infinity-shaped reference.
A detailed literature review can be found in the subsequent section. Existing works typically rely on the following assumptions: 1) The inner-loop controller is known and accessible to the user; 2) The inner-loop controller is known but not accessible to the user; 3) The inner-loop controller has perfect tracking for outer-loop commands. Our goal is to minimize reliance on knowledge of the dynamic model and the inner-loop controller, while still guaranteeing tracking and safe control using only an outer-loop kinematic controller.
As shown in Fig. 2, existing robotic systems typically send desired position ($q^{*}$) or velocity ($\dot{q}^{*}$) commands to the inner-loop controller. Our solution, inserted between them (gray box), comprises a disturbance-rejection tracking controller, an extended state observer (ESO), and an extended state observer-based robust quadratic programming (ESOR-QP) module. The ESO estimates the total disturbance $\hat{f}$ using the joint angle $q$ and the safe kinematic control $q_{d,\text{safe}}$ or $\dot{q}_{d,\text{safe}}$. This estimate is compensated in the tracking controller to produce nominal commands $q_{d}$ or $\dot{q}_{d}$, which are monitored by the ESOR-QP to enforce robust safety constraints.
The contributions of our work are summarized as follows:
  • We innovatively formulate the kinematic control of closed-architecture robotic manipulator systems as a disturbance rejection problem, requiring only a practical assumption on the structure, Jacobian, and nominal inertia matrix.
  • We develop a robust control barrier function (CBF) framework based on an extended state observer, enabling safe control of general nonlinear affine systems with high relative degrees under disturbances. To the best of our knowledge, this is the first work achieving safe kinematic control of closed-architecture robots under largely unknown uncertainties.
  • As shown in Fig. 2, our method is implemented as a standalone add-on module that enhances existing robotic manipulators without modifying the inner-loop controller. Real-time control at 1 kHz is demonstrated on a PUMA 500 manipulator.
Figure 2: The proposed framework in this paper, is marked as the Outer-loop Controller. The total estimated disturbance $\hat{f}$ is used for both disturbance compensation and safe control using ESOR-QP.
The rest of the paper is organized as follows: Section II reviews the related work. Section III introduces the preliminaries and problem formulation. Section IV presents the proposed framework. Hardware experiments are reported in Section V. Concluding remarks are in Section VI.
II Related Work
An overview of the comparison between our work and the most related works is presented in Table I.
II-A Adaptive Control for Closed Architecture
Learning-based Adaptive Control: [2, 3, 4] employ neural network–based outer-loop adaptation to approximate the kinematics, dynamics, and inner-loop controller. As summarized in Table I, they make the weakest assumption about the inner loop, identical to ours, where $\Psi(\cdot)$ denotes an unknown structure possibly containing proportional, integral, or state-dependent terms. However, they rely on numerous tuning parameters and limited interpretability.
Conventional Adaptive Control: [1] introduces dynamic modularity with an adaptive controller under a moderate assumption of a known PI/PID inner-loop structure.
These studies address tracking control only, without safety considerations. In our experiments, we compare tracking performance with the learning-based adaptive control due to the shared inner-loop assumption.
TABLE I: Overview of our work’s position in the literature
Methods Tracking Safety
Assumption of
Inner-loop Controller
Ref.
Learning-based
Adaptive Control
Weak
$\tau=-K_{d}(\dot{q}-\dot{q}_{d})+\Psi(q,\dot{q},q_{d},\int_{0}^{t}q\,dt,\int_{0}^{t}q_{d}\,dt)$
[2, 3, 4]
Conventional
Adaptive Control
Moderate, known PID/PI structure
[1]
Reduced-order CBF -
Strong, known PD structure,
perfect tracking or
bounded error
[6, 7, 8]
Robust CBF - - [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]
Ours
Weak
$\tau=-K_{d}(\dot{q}-\dot{q}_{d})+\Psi(q,\dot{q},q_{d},\int_{0}^{t}q\,dt,\int_{0}^{t}q_{d}\,dt)$
  • The two dashes for CBFs’ tracking indicate that we only compare our safety performance with them, as their nominal tracking controllers vary significantly across papers. The third dash for the robust CBF indicates that those works do not consider two-loop controllers.
II-B Reduced-order Model-based CBF
Many CBF-based safety approaches rely on high-fidelity dynamic models with torque control, yet obtaining such models remains challenging. Reduced-order or kinematic CBFs have been introduced to ensure safety with lower-fidelity models. For instance, [6] guarantees safety via safe velocity control but assumes a perfect inner-loop controller capable of instant tracking. Similarly, [7, 8] employ kinetic energy–based CBFs that offer partial robustness to model uncertainty but still depend on numerous dynamic parameters and a known PD-type inner loop. In contrast, our work addresses closed-architecture robots with minimal knowledge of the inner-loop controller and significant model uncertainty.
II-C Robust CBF
Robust CBFs have recently been developed to handle model uncertainties and external disturbances [9, 10, 11, 12, 13, 14, 15, 17, 19, 20, 21, 25, 18, 16, 23, 24, 22]. Among them, disturbance observer–based robust CBFs (DOB-CBFs) [25, 18, 16, 23, 24, 22, 15, 17, 19, 20, 21] enable active disturbance estimation without requiring worst-case bounds. Unlike these methods where estimation error bounds are absent or loosely derived via Lyapunov analysis, our approach obtains them directly from the observer error dynamics, resulting in tighter and less conservative bounds. Experimental comparisons with state-of-the-art methods (e.g., [16]) further demonstrate this advantage. In contrast to our conference version [25] that only considered relative-degree-one systems, this work provides a generalized framework with rigorous proofs guaranteeing the forward invariance of the safety set.
III Preliminary and Problem Formulation
III-A System Dynamics
We consider a general robotic system with an Euler–Lagrange formulation incorporating a generalized external disturbance force $F_{ext}$:
$\begin{cases}{M}\ddot{q}+{C}\dot{q}+{G}+F_{r}=\tau+{J}^{T}F_{ext}\\ \hfil y=\zeta(q)\end{cases},$ (1)
where $q,\dot{q},\ddot{q}\in\mathbb{R}^{n}$ are the joint position, velocity, and acceleration, respectively. ${M}\in\mathbb{R}^{n\times n}$ is the symmetric positive-definite inertia matrix; ${C}\in\mathbb{R}^{n\times n}$ denotes centrifugal and Coriolis term; ${G}\in\mathbb{R}^{n}$ is the gravity term; $F_{r}\in\mathbb{R}^{n}$ accounts for the friction torque; $\tau\in\mathbb{R}^{n}$ represents the torque (provided by the inner-loop controller); ${J}\in\mathbb{R}^{6\times n}$ is the Jacobian matrix; and $F_{ext}\in\mathbb{R}^{6}$ represents a generalized external disturbance force. The output $y\in\mathbb{R}^{m}$, and $\zeta:\mathbb{R}^{n}\to\ \mathbb{R}^{m}$ is the mapping from joint space to task space.
The most common inner-loop controller for robot manipulators typically consists of a PID controller or PD controller, often combined with dynamic compensation, such as gravity. For stability, most inner-loop controllers include a derivative term for damping, $(\dot{q}-\dot{q}_{d})$, where $\dot{q}_{d}$ is a joint velocity command.
Assumption 1
We assume the inner-loop controller has the following structure:
$\tau=-K_{d}(\dot{q}-\dot{q}_{d})+\Psi(q,\dot{q},q_{d},\int_{0}^{t}q\,dt,\int_{0}^{t}q_{d}\,dt),$ (2)
with $K_{d}$ a proportional control gain, and $\Psi(\cdot)$ is a general, unknown function that may include proportional, integral, derivative, and dynamic compensation terms. To the best of our knowledge, this assumption is the most general in the literature (see Table I), requiring minimal knowledge of the inner-loop controller, similar to [2], and more general than [1].
Assumption 2
The kinematics of the system are assumed to be known for control in task space.
The system dynamic model combining (1) and (2) is:
$\begin{split}\ddot{q}&={M}^{-1}\Bigl(-{C}\dot{q}-{G}-F_{r}+{J}^{T}F_{ext}-{K}_{d}\dot{q}+\Psi(\cdot)+{K}_{d}\dot{q}_{d}\Bigr)\end{split}.$ (3)
We have removed torque control, as it is not accessible in a closed architecture. $\dot{q}_{d}$ is the kinematic control to be manipulated in the outer-loop layer, and its control law will be designed. Another reason for choosing $\dot{q}_{d}$ instead of $q_{d}$ as the control input is to avoid differentiation-induced noise.
III-B Robot Kinematic Control with Known Model
Neglecting $F_{r}$ and $F_{ext}$ in (1), the exact system dynamics is used to design the control law:
$\begin{split}\dot{q}_{d}&={K}_{d}^{-1}\Bigl({M}u_{0}+{C}\dot{q}+{G}+{K}_{d}\dot{q}-\Psi(\cdot)\Bigr)\end{split},$ (4)
where the position control input $u_{0}$ is defined as follows:
$\begin{split}u_{0}&=\ddot{q}^{\star}+k_{p}(q^{\star}-q)+k_{d}(\dot{q}^{\star}-\dot{q})\\ \end{split},$ (5)
where $\ddot{q}^{\star},\dot{q}^{\star},q^{\star}\in\mathbb{R}^{n}$ represent the desired acceleration, velocity, and position of a reference joint trajectory generated by an upstream planner. $k_{p}\in\mathbb{R}^{n\times n}$ and $k_{d}\in\mathbb{R}^{n\times n}$ are the proportional and derivative gain matrices, respectively. $\ddot{q}^{\star}$ serves only as a feedforward term, while $u_{0}$ is a position controller. By substituting (4) and (5) into (3), we get:
$\ddot{q}=\ddot{q}^{\star}+k_{p}(q^{\star}-q)+k_{d}(\dot{q}^{\star}-\dot{q}).$ (6)
Proper $k_{p}$ and $k_{d}$ make q{q} converge to ${q}^{\star}$ asymptotically.
III-C Control Barrier Functions
The system (3) can be rewritten as a nonlinear control-affine system when ignoring the disturbance vector $F_{ext}$:
$\dot{x}=\psi(x)+g(x)u,$ (7)
where $x\triangleq[q,\dot{q}]^{T}\in\mathcal{X}\subset\mathbb{R}^{2n}$, $\psi:\mathbb{R}^{2n}\rightarrow\mathbb{R}^{2n}$ and $g:\mathbb{R}^{2n}\rightarrow\mathbb{R}^{2n\times n}$ are Lipschitz continuous, and $u\triangleq\dot{q}_{d}\in\mathcal{U}\subset\mathbb{R}^{n}$ is a control input vector.
The safety of system (7) can be guaranteed using a safety set. A set $\mathcal{C}$ is considered to be a safety set if it is forward invariant in the state space $\mathcal{X}$, i.e., for system (7) if solutions for some $u\in\mathcal{U}$ starting at any initial safe condition $x(0)\in\mathcal{C}$ satisfy $x(t)\in\mathcal{C}$, $\forall t\geq 0$. The safety set $\mathcal{C}$ is defined as a 0-superlevel set of a continuously-differential function $h(x):\mathbb{R}^{2n}\rightarrow\mathbb{R}$ as:
$\mathcal{C}=\{x\in\mathbb{R}^{2n}~|~h(x)\geq 0\}.$ (8)
The function $h$ is used to synthesize a controller with safety guarantees via a CBF.
Definition 1
(Exponential CBF (ECBF) [26, 27]) Consider system (7) with relative degree $r$ for an $r$-times continuously differentiable function $h$, i.e., $L_{g}L_{\psi}h(x)=\cdots=L_{g}L_{\psi}^{r-2}h(x)=0$ and $L_{g}L_{\psi}^{r-1}h(x)\neq 0,\forall x\in\mathcal{C}$. $h(x)$ is an ECBF if there exists a row vector $K_{a}\in\mathbb{R}^{r}$ satisfying $\forall x\in\mathcal{C}$
$\sup\limits_{u\in\mathcal{U}}\big(L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u\big)\geq-K_{a}\eta_{b}(x),$ (9)
where $\eta_{b}(x)=[h(x),\dot{h}(x),\cdots,h^{(r-1)}(x)]^{T}$, $K_{a}=[k_{1},\cdots,k_{r}]$, and the values of $k_{1},\cdots,k_{r}$ satisfy specific properties given in [26, 27].
To guarantee safety, the control problem is formulated as a quadratic program (QP) with a CBF as a hard constraint [28]:
$\begin{split}u^{*}(x)&=\operatorname*{arg\,min}_{u\in U}\|u-k(x)\|^{2}\\ \text{s.t.}\quad L_{\psi}^{r}h(x)&+L_{g}L_{\psi}^{r-1}h(x)u\geq-K_{a}\eta_{b}(x),\end{split}$ (10)
where $k(x)$ is a nominal control law.
Research Objective: Given the system in (3), the nominal dynamics with limited knowledge of the system and inner-loop controller are defined as
$\ddot{q}=\bar{M}^{-1}\!\left(-\bar{C}\dot{q}-\bar{G}+\bar{K}_{d}\dot{q}_{d}\right),$ (11)
where $\bar{M}$, $\bar{C}$, and $\bar{G}$ are nominal parameters, and $\bar{K}_{d}$ is the nominal inner-loop gain. The goal is to design a $\dot{q}_{d}$ controller that achieves accurate trajectory tracking and task-space safety (e.g., collision avoidance) under an unknown inner-loop controller, uncertain dynamics, and external disturbances.
Remark 1
$\bar{C}$ and $\bar{G}$ can be a zero matrix and a zero vector, respectively, with minimal knowledge of the system. We use $\bar{C}$ and $\bar{G}$ here to make the formulation more general. The model (11) can be further reduced to:
$\ddot{q}=\bar{M}^{-1}\bar{K}_{d}\dot{q}_{d}.$ (12)
We use the nominal model (11) for general mathematical derivation, whereas the reduced nominal model (12) is employed in hardware experiments to evaluate the tracking performance and safety under limited information.
IV Proposed Framework
IV-A Robot Kinematic Control with Nominal Model
The control law in (4) is ideal, assuming known dynamics and an inner-loop controller. To handle uncertainties, we propose the following practical law based on nominal dynamics:
$\begin{split}\dot{q}_{d}&=\bar{K}_{d}^{-1}\Bigl(\bar{M}u_{0}+\bar{C}\dot{q}+\bar{G}\Bigr).\end{split}$ (13)
The following part shows the dynamics of the close-loop system with our nominal control law (13). We define the parametric discrepancies as:
$\begin{split}\Delta K_{d}&\triangleq K_{d}\bar{K}_{d}^{-1},\\ \Delta{M}&\triangleq\Delta K_{d}\bar{M}-{M},\\ \Delta{C}&\triangleq\Delta K_{d}\bar{C}-{C},\\ \Delta{G}&\triangleq\Delta K_{d}\bar{G}-{G}.\end{split}$ (14)
By rearranging (3), we get:
[…]
Substituting (13) and (14) into (15), the right-hand side of (15) becomes:
[…]
The left-hand side of (15) becomes:
$\begin{split}&\Bigl(\Delta K_{d}\bar{M}-\Delta{M}\Bigr)\ddot{q}+\Bigl(\Delta K_{d}\bar{C}-\Delta{C}\Bigr)\dot{q}\\ &+\Bigl(\Delta K_{d}\bar{G}-\Delta{G}\Bigr)+F_{r}.\end{split}$ (17)
The resulting close-loop system is:
$\begin{split}\ddot{q}=\ddot{q}^{\star}+k_{p}(q^{\star}-q)+k_{d}(\dot{q}^{\star}-\dot{q})+f.\end{split}$ (18)
where
$\begin{split}f=(\Delta K_{d}\bar{M})^{-1}\Bigl(&\Delta{M}\ddot{q}+\Delta{C}\dot{q}+\Delta{G}-F_{r}\\ &+\Psi(\cdot)+{J}^{T}F_{ext}-K_{d}\dot{q}\Bigr).\end{split}$ (19)
Let $f$ denote the total disturbance, including internal uncertainties and external disturbances. Compared with (6), $f$ degrades system performance. To counter this, an ESO [29] is employed to estimate $f$ for disturbance rejection.
IV-B Extended State Observer Design
For the $i$-th joint of the robotic manipulator (11), we have the following second-order subsystem:
$\begin{cases}\begin{array}[]{l}\dot{x}_{1i}=x_{2i}\\ \dot{x}_{2i}=F_{i}(x)+G_{i}(x)\dot{q}_{di}+f_{i},\end{array}\end{cases}$ (20)
where $x_{1i}=q_{i}$; $x_{2i}=\dot{q}_{i}$; $F_{i}(x)$, $G_{i}(x)$, and $f_{i}$ are the $i$-th row of $F=\bar{M}^{-1}(-\bar{C}\dot{q}-\bar{G})\in\mathbb{R}^{n}$, $G=\bar{M}^{-1}\bar{K}_{d}\in\mathbb{R}^{n}$, and $f$, respectively.
By treating $f_{i}$ as an extended state, the augmented system can be written as:
$\begin{cases}\begin{array}[]{l}\dot{x}_{1i}=x_{2i}\\ \dot{x}_{2i}=F_{i}(x)+G_{i}(x)\dot{q}_{di}+x_{3i}\\ \dot{x}_{3i}=\dot{f}_{i}.\end{array}\end{cases}$ (21)
In many robotic systems, both joint position $q$ and velocity $\dot{q}$ are measurable, enabling the design of second- or third-order ESOs based on (21). With velocity measurements available, a second-order ESO is proposed for the $i$-th joint to estimate $f_{i}$; otherwise, it can be readily extended to a third-order form.
The second-order ESO can be designed as:
$\begin{cases}\begin{array}[]{l}\dot{\hat{x}}_{2i}=F_{i}(x)+G_{i}(x)\dot{q}_{di}+\hat{x}_{3i}+\beta_{1i}(x_{2i}-\hat{x}_{2i})\\ \dot{\hat{x}}_{3i}=\beta_{2i}(x_{2i}-\hat{x}_{2i}),\end{array}\end{cases}$ (22)
where $\hat{x}_{2i}$ and $\hat{x}_{3i}$ are the estimated velocity for $\dot{q}_{i}$ and estimated disturbance for $f_{i}$, respectively. $\beta_{1i}$ and $\beta_{2i}$ are observer gains. The estimation error dynamic is derived from (21) and (22):
$\dot{\vartheta}_{i}=A_{i}\vartheta_{i}+E_{i}\dot{f}_{i},\\$ (23)
where $\vartheta_{i}=\begin{bmatrix}x_{2i}-\hat{x}_{2i}\\ x_{3i}-\hat{x}_{3i}\\ \end{bmatrix}$, $A_{i}=\begin{bmatrix}-\beta_{1i}&1\\ -\beta_{2i}&0\\ \end{bmatrix}$, $E_{i}=\begin{bmatrix}0\\ 1\end{bmatrix}$.
The observer gains are chosen such that the eigenvalues of the matrix $A_{i}$ are placed at $\omega_{o_{i}}$, where $\omega_{o_{i}}$ is the only tunable value known as the observer bandwidth [30].
IV-C Controller Design
$\hat{f}$ can be used to develop a disturbance rejection controller that complements the control law (13):
[…]
where $\dot{q}_{d_{0}}$ is the nominal control from (13), $\dot{q}_{d_{f}}$ rejects disturbance, and they are summed to get $\dot{q}_{d}$.
IV-D Controller Stability Analysis
Assumption 3
There exists a positive known constant $l_{f}$ such that for any $x\in\mathcal{X}$, $u\in\mathcal{U}$, and $t\geq 0$, the following inequality holds:
$\left|\dfrac{\partial f_{i}(t,x,u)}{\partial t}\right|\leq l_{f}.$ (25)
Assumption 3 implies that $f$ is locally Lipschitz continuous with respect to time, and $\dot{f}$ is bounded.
Theorem 1
Given the robotic system (3) with kinematic control input, the ESO (22) with appropriate observer bandwidth $\omega_{o_{i}}$ for each joint, and the nominal control law (24) with appropriate tuned $k_{p}$ and $k_{d}$, we have the following properties: (i): the closed-loop system of (3) is bounded stable if each $\omega_{o_{i}}$ is finite; (ii): the closed-loop system of (3) is asymptotically stable if each $\omega_{o_{i}}$ approaches infinity.
Proof 1
Since $A_{i}$ is Hurwitz, we have the following decomposition [31, Chapter 1.3]:
$e^{A_{i}t}=P_{i}e^{\Lambda_{i}t}P_{i}^{-1},$ (26)
where $P_{i}$ is a matrix whose columns are the eigenvectors of $A_{i}$, and $\Lambda_{i}$} is a diagonal matrix with the corresponding eigenvalues in the diagonal elements.
Therefore, there exists a constant $c_{1_{i}}$ such that [31, Chapter 1.9]
$\left\lVert e^{A_{i}t}\right\rVert\leq\left\lVert P_{i}\right\rVert\left\lVert e^{\Lambda_{i}t}\right\rVert\left\lVert P_{i}^{-1}\right\rVert=c_{1_{i}}\left\lVert e^{\Lambda_{i}t}\right\rVert\leq c_{1_{i}}\left\lVert e^{a_{i}t}\right\rVert,$ (27)
where $a_{i}$ is the $\lambda_{max}(\Lambda_{i})$, the maximum eigenvalue of $\Lambda_{i}$, which, in our case, is $\omega_{o_{i}}$.
The solution to the error dynamics in (23) is given by:
$\vartheta_{i}=e^{A_{i}(t-t_{0})}\vartheta_{i}(t_{0})+\int_{t_{0}}^{t}e^{A_{i}(t-\tau)}E\dot{f}_{i}(\tau)\,d\tau,$ (28)
where $t_{0}$ is the initial time. Substituting (27) into (28), we obtain:
[…]
Since $\dot{f}_{i}$ is bounded, $e^{-\omega_{o_{i}}(t-t_{0})}$ decays exponentially, the estimation error $\left\lVert\vartheta_{i}\right\rVert\ \leq\sigma_{i}$ is bounded. With $\omega_{o_{i}}$ approaching infinity, $\sigma_{i}$ approaches 0.
Thus, the disturbance estimation error satisfies:
$\left\lVert\tilde{f}\right\rVert=\left\lVert f-\hat{f}\right\rVert\leq\sigma.$ (30)
For the $i$-th joint subsystem, substituting the control law (24) into (3) results in the following $i$-th closed-loop subsystem dynamics:
$\ddot{q}_{i}=\ddot{q}_{i}^{\star}+k_{p_{i}}(q_{i}^{\star}-q_{i})+k_{d_{i}}(\dot{q}_{i}^{\star}-\dot{q}_{i})+\tilde{f}_{i},$ (31)
where $\tilde{f}_{i}=f_{i}-\hat{f}_{i}$ represents the disturbance estimation error. The error dynamic can be expressed as:
$\dot{\epsilon}=H\epsilon+\xi,$ (32)
where $\epsilon=\begin{bmatrix}q_{i}-q_{i}^{\star}\\ \dot{q}_{i}-\dot{q}_{i}^{\star}\\ \end{bmatrix}$, $H=\begin{bmatrix}0&1\\ -k_{p_{i}}&-k_{d_{i}}\\ \end{bmatrix}$, and $\xi=\begin{bmatrix}0\\ \tilde{f}_{i}\\ \end{bmatrix}$.
The matrix $H$ can be designed to be Hurwitz. Choose $V=\frac{1}{2}\epsilon^{T}Q\epsilon$ as the Lyapunov function candidate. Its derivative along (32) is given by:
$\dot{V}=\tfrac{1}{2}(\dot{\epsilon}^{T}Q\epsilon+\epsilon^{T}Q\dot{\epsilon})=-\|\epsilon\|^{2}+\xi^{T}Q\epsilon.$ (33)
By applying the Cauchy–Schwarz inequality, we obtain:
[…]
Given that $Q$ is positive definite, $\left\lVert Q\right\rVert\leq\lambda_{max}(Q)$, and the $\xi$ is bounded by $\sigma$, the inequality (34) becomes:
$\begin{array}[]{ll}\dot{V}&\leq-\left\lVert\epsilon\right\rVert^{2}+\sigma\lambda_{max}(Q)\left\lVert\epsilon\right\rVert.\end{array}$ (35)
Next, using the inequality:
$\begin{array}[]{ll}\sigma\lambda_{max}(Q)\left\lVert\epsilon\right\rVert\leq\frac{1}{2}\left\lVert\epsilon\right\rVert^{2}+\frac{1}{2}\sigma^{2}\lambda^{2}_{max}(Q),\end{array}$ (36)
(35) can be rewritten as:
$\begin{array}[]{ll}\dot{V}&\leq-\frac{1}{2}\left\lVert\epsilon\right\rVert^{2}+\frac{1}{2}\sigma^{2}\lambda^{2}_{max}(Q)\\ &\leq-\frac{V}{\lambda_{max}(Q)}+\frac{1}{2}\sigma^{2}\lambda^{2}_{max}(Q).\end{array}$ (37)
Solving the inequality (37) yields:
$\begin{array}[]{ll}{V}&\leq(v(0)-\frac{\kappa}{\eta})e^{-\eta t}+\frac{\kappa}{\eta},\end{array}$ (38)
where $\eta=\frac{1}{\lambda_{max}(Q)}$ and $\kappa=\frac{1}{2}\sigma^{2}\lambda^{2}_{max}(Q)$. 1) It can be concluded from (38) that the states of the closed-loop system are bounded. 2) Moreover, as $\omega_{o}$ approaches infinity, $\sigma$ approaches zero, reducing the disturbance estimation error. Consequently, the inequality (35) is simplified as $\dot{V}\leq-\eta V$, which implies that the system is asymptotically stable.
IV-E Robust CBF for Safe Control
By adding the disturbance term, system (11) becomes:
$\dot{x}=\psi(x)+g(x)u+\bar{g}f,$ (39)
where $f=[f_{1},\cdots,f_{n}]^{T}\in\mathcal{F}\subset\mathbb{R}^{n}$ is a total disturbance vector in each input channel, $\bar{g}=[\mathbf{0}_{n\times n},\mathbf{I}_{n\times n}]^{T}$, $\mathbf{0}_{n\times n}$ and $\mathbf{I}_{n\times n}$ are $n\times n$ zero and identity matrices, respectively. Note that $u$ and $f$ are in the same channel, i.e., the $n\times n$ matrix of the upper part of $g(x)$ is a zero matrix and each row in the lower part of $g(x)$ includes an independent input, so $f$ is a matched disturbance vector.
To define safety sets for system (39), we consider an $r$-times continuously differentiable function $h(x)$, where the relative degrees of $h(x)$ with respect to both $u$ and $f$ are $r$, given that $f$ is matched. A series of functions are defined as follows:
$\begin{array}[]{ll}h_{0}(x)=&h(x),\\ h_{1}(x)=&\dot{h}_{0}(x)+\gamma_{1}h_{0}(x),\\ &\vdots\\ h_{r}(x)=&\dot{h}_{r-1}(x)+\gamma_{r}h_{r-1}(x),\end{array}$ (40)
where $\gamma_{1},\cdots,\gamma_{r}$ are positive constants. The corresponding series of safety sets are
$\begin{array}[]{ll}\mathcal{C}_{0}=&\{x\in\mathbb{R}^{2n}:h_{0}(x)\geq 0\},\\ \mathcal{C}_{1}=&\{x\in\mathbb{R}^{2n}:h_{1}(x)\geq 0\},\\ &\vdots\\ \mathcal{C}_{r}=&\{x\in\mathbb{R}^{2n}:h_{r}(x)\geq 0\}.\end{array}$ (41)
Theorem 2
For system (39), the relative degrees of $h(x)$ with respect to both the input $u$ and the disturbance $f$ are $r$, and $f$ is known. If the initial states satisfy $x(0)\in\mathcal{C}_{s}=\cap_{j=0}^{r}\mathcal{C}_{j}$, then any Lipschitz continuous controller $u(x)\in K_{\text{cbf}}(t,x,f)$ renders the set $\mathcal{C}_{s}$ forward invariant for system (39), where
$\begin{array}[]{r@{}l}K_{\text{cbf}}(t,x,f)\triangleq&\{u\in\mathcal{U}:L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u\\ &+L_{\bar{g}}L_{\psi}^{r-1}h(x)f+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\geq 0\},\end{array}$ (42)
where $h^{(j)}(x)$ is the $j$th time derivative of $h(x)$, and $k_{j},j=0,\cdots,r-1$, are the coefficients of polynomial $s^{r}+k_{r-1}s^{r-1}+\cdots+k_{0}$ with roots at $\gamma_{1},\cdots,-\gamma_{r}$.
Proof 2
From (40), $h_{r}$ can be written as
$\begin{array}[]{r@{}l}h_{r}(x)=&(s+\gamma_{r})(s+\gamma_{r-1})\cdots(s+\gamma_{1})h(x)\\ =&s^{r}h(x)+k_{r-1}s^{r-1}h(x)+\cdots+k_{0}h(x),\end{array}$ (43)
where $s\triangleq\frac{d}{dt}$. Since the control input $u$ and disturbance $f$ both have the same relative degree of $r$, $u$ and $f$ do not explicitly show in $h^{(j)}(x),0\leq j\leq r-1$ until
$h^{(r)}(x)=L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)f.$ (44)
By substituting (44) into (43) and comparing it with (42), the control values in $K_{\text{cbf}}(t,x,f)$ guarantee $h_{r}(x)\geq 0$, $\forall t>0$. From the last equation in (40), we have $\dot{h}_{r-1}(x)+\gamma_{r}h_{r-1}(x)\geq 0$. Then $\dot{h}_{r-1}(x)\geq 0$ for any $x\in\partial\mathcal{C}_{r-1}$. According to Nagumo’s theorem [32], since $x(0)\in\mathcal{C}_{r-1}$, we have $h_{r-1}(x)\geq 0$, $\forall t>0$. Iteratively, $\dot{h}_{0}(x)+\gamma_{1}h_{0}(x)\geq 0$, and we have $h_{0}(x)\geq 0$, $\forall t>0$, as $x(0)\in\mathcal{C}_{0}$. Therefore, the set $\mathcal{C}_{s}$ is forward invariant for system (39). This completes the proof of Theorem 2.
However, the disturbance vector $f$ is not available in practice to enforce the inequality in (42). The estimated disturbance vector $\hat{f}$ obtained in Subsection IV-B is used to devise our robust CBF. Then, (42) can be reformulated as:
$\begin{array}[]{r@{}l}K_{\text{cbf}}(t,x,f,\hat{f})\triangleq&\{u\in\mathcal{U}:L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u\\ &+L_{\bar{g}}L_{\psi}^{r-1}h(x)\hat{f}-L_{\bar{g}}L_{\psi}^{r-1}h(x)(\hat{f}-f)\\ &+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\geq 0\}.\end{array}$ (45)
Since the disturbance estimation error $\hat{f}-f$ is unknown in (45), the following assumption is made for the disturbance estimation error bound of ESO.
The estimation error of $\hat{f}_{i}$ is as follows [29]:
$f_{i}(k)-\hat{f}_{i}(k)=p(k)*\Delta f_{i}(k),$ (46)
where * represents convolution and
$\footnotesize\begin{array}[]{l}p(k)=\\ \begin{cases}1&1\leq k\leq r_{i}+1\\ \sum\limits_{\iota=1}^{r_{i}+1}\dfrac{1}{(\iota-1)!}\left(\prod\limits_{j=-\iota+1}^{-1}(k+j)\right)(1-\omega_{o_{i}})^{\iota-1}\omega_{o_{i}}^{k-\iota}&k\geq r_{i}+2.\end{cases}\end{array}$ (47)
For simplicity, let $\prod_{j=0}^{-1}(k+j)=1$. Note that (46) is the disturbance estimation error formulated in the discrete-time domain. $f_{i}(k)=f_{i}(kT_{s})$, $\hat{f}_{i}(k)=\hat{f}_{i}(kT_{s})$, $\Delta f_{i}(k)=f_{i}((k+1)T_{s})-f_{i}(kT_{s})$, $T_{s}$ is the sample time used in calculating the error bound, and $r_{i}$ and $\omega_{o_{i}}$ are the relative degree of $\dot{q}_{i}$ (output measurement) with respect to $f_{i}$ (disturbance as input) in the $i$-th joint subsystem and the poles of the $i$-th ESO in the discrete-time domain, respectively. Readers are referred to our previous work [29] for a complete proof of this error bound.
From the definition of derivative, Assumption 3, and (46), the disturbance estimation error bound of ESO is
$|f_{i}(k)-\hat{f}_{i}(k)|\leq\Gamma_{i}(\omega_{o_{i}},T_{s})=\left(\sum\limits_{k=1}^{\infty}p(k)\right)l_{f}T_{s}.$ (48)
As mentioned in [25], although (48) is derived in the discrete-time domain, the disturbance estimation error bound obtained from it is still the same as that from the continuous-time domain. Since we need to utilize the disturbance estimation error bound in the continuous-time domain, $T_{s}$ should be small, such as 0.1 ms, in our experiment setting. $\omega_{o_{i}}$ in the discrete-time domain needs to be converted from the continuous-time domain through Z-transform.
Theorem 3
Given the system (39) and the ESO in (22) under Assumption 3, any controller $u(x)\in K_{\text{rcbf}}$ renders the set $\mathcal{C}_{s}$ forward invariant for system (39), where
$\begin{array}[]{r@{}l}K_{\text{rcbf}}(t,x,f,\hat{f})\triangleq&\{u\in\mathcal{U}:L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u\\ &+L_{\bar{g}}L_{\psi}^{r-1}h(x)\hat{f}-|L_{\bar{g}}L_{\psi}^{r-1}h(x)|\Gamma(\omega_{o},T_{s})\\ &+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\geq 0\},\end{array}$ (49)
$|L_{\bar{g}}L_{\psi}^{r-1}h(x)|$ denotes the absolute value of each element, and $\Gamma(\omega_{o},T_{s})=[\Gamma_{1}(\omega_{o_{1}},T_{s}),\cdots,\Gamma_{n}(\omega_{o_{n}},T_{s})]^{T}$.
Proof 3
From Theorem 2, we need to prove $h_{r}(x)\geq 0$, $\forall t>0$. Substituting (44) and (48) into (43) yields
$\begin{array}[]{r@{}l}h_{r}(x)=&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)f\\ &+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\\ =&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)\hat{f}\\ &-L_{\bar{g}}L_{\psi}^{r-1}h(x)(\hat{f}-f)+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\\ \geq&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)\hat{f}\\ &-|L_{\bar{g}}L_{\psi}^{r-1}h(x)|\Gamma(\omega_{o},T_{s})+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\geq 0.\end{array}$ (50)
Then, we have $h_{r}(x)\geq 0$, $\forall t>0$ as $x(0)\in\mathcal{C}_{r}$. Using the same procedure in Theorem 2, the set $\mathcal{C}_{s}$ is forward invariant for system (39). This completes the proof of Theorem 3.
Remark 2
The key difference between our theoretical results in this work and our preliminary version [25] is that we have extended the robust CBF’s safety guarantee from a relative degree of one to a general case of higher order CBF.
The safety specification requires the end-effector to maintain a non-negative distance from a virtual wall. The safety function is defined as $h(x)=y-y_{0}$, where $y$ and $y_{0}$ denote the positions of the end-effector and the virtual wall, respectively. It can be verified that $h(x)$ has a relative degree of two [26]:
[…]
where a positive constant $\gamma_{1}=\gamma_{2}=\gamma$ is chosen,
$\dot{y}=\left(\frac{\partial\zeta}{\partial q}\right)\dot{q},~\text{and}~\ddot{y}=\frac{d}{dt}\left(\frac{\partial\zeta}{\partial q}\right)\dot{q}+\frac{\partial\zeta}{\partial q}\ddot{q}.$
Combining equations (11), (19), and (51) yields
$\begin{array}[]{ll}h_{2}=&2\gamma\dot{y}+\gamma^{2}(y-y_{0})+\frac{d}{dt}\left(\frac{\partial\zeta}{\partial q}\right)\dot{q}\\ &+\frac{\partial\zeta}{\partial q}\Bigl(-\bar{M}^{-1}(\bar{C}\dot{q}+\bar{G}-\bar{K}_{d}\dot{q}_{d})+f)\Bigr).\end{array}$ (52)
To simultaneously achieve tracking and safe control, the nominal controller for $\dot{q}_{d}$, designed in (24), should be minimally intervened. Thus, the following QP-CBF is constructed:
$\dot{q}_{d_{\text{safe}}}=$ $\operatorname*{arg\,min}_{u\in\mathcal{U}}\|u-\dot{q}_{d}\|^{2}$ (53)
s.t. $h_{i}(x(0))>0,~i=0,1,2$
${L_{\psi}^{2}h(x)}+L_{g}L_{\psi}h(x)u$
$+L_{\bar{g}}L_{\psi}h(x)\hat{f}-{|L_{\bar{g}}L_{\psi}h(x)|}\Gamma(\omega_{o},T_{s})$
$+2\gamma\dot{h}(x)+\gamma^{2}h(x)\geq 0.$
V Hardware Experiments
We use the PUMA 500 robot for hardware experiments, as shown in Fig. 3. It has three main joints and a spherical wrist, providing six degrees of freedom. $q_{1}$, $q_{2}$, and $q_{3}$ correspond to the waist, shoulder, and elbow joints, while $q_{4}$, $q_{5}$, and $q_{6}$ represent the wrist rotation, bend, and flange angles. Since our focus is task-space tracking (i.e., wrist center position), only the first three joints are used. Each joint is driven by a DC motor powered by an Adept Technology MV-19 Power Chassis. The control program is built in MATLAB Simulink, which generates C code and an SDF file, and runs via a dSPACE 1103 for control and data acquisition.
Figure 3: PUMA 500 robot used as a testbed in our research.
The inner-loop controller is configured as $\tau=-K_{p}({q}-{q}_{d})-K_{d}(\dot{q}-\dot{q}_{d})$, where $K_{p}=\mathrm{diag}([18,144,9])$ and $K_{d}=\mathrm{diag}([18,12,9])$. The gains are non-round values due to the actuator driver’s voltage-to-torque conversion ratio. The structure and true parameters remain unknown to our outer-loop controller. The nominal second-order system model (12) is used for the controller. The reference trajectory for joints $q_{1}$, $q_{2}$, and $q_{3}$ is $\begin{bmatrix}0.5\sin(t)&0.25\sin(2t)+0.534&0.25\sin(2t)+\frac{\pi}{2}\end{bmatrix}^{T}$. The tracking performance of the nominal inner-loop controller is illustrated in Fig. 4. Specifically, Fig. 4a depicts the trajectory tracking performance in the joint space, while Fig. 4b shows the trajectory tracking performance in the Cartesian space. The results indicate that the inner-loop controller exhibits poor tracking performance. We will demonstrate the improvement achieved by adding our approach to the tracking and safe control tasks.
We use the nominal model (12) throughout the experiments. A key parameter in the controller design is the control gain in (24), $b_{0}\triangleq\bar{M}^{-1}\bar{K}_{d}$, which depends on the nominal inertia matrix $\bar{M}$ and the inner-loop proportional gain. The nominal inertia $\bar{M}$ is adopted from [33]. Following a common ESO tuning rule, we tune each joint individually by initially setting $b_{0}$ large to ensure stability and then gradually reducing it to improve tracking accuracy. The final nominal control gain is $\mathrm{diag}([20,40,10])$. The observer and control bandwidths are set to $80~\mathrm{rad/s}$ and $10~\mathrm{rad/s}$, respectively, for each joint. We refer readers to [34] for additional instructions and theoretical support on tuning nominal control gain.
(a) Joint space performance
(b) Task space performance
Figure 4: Poor inner-loop PD controller’s performance for $10$ s. (a) Trajectory tracking performance in the joint space. (b) Trajectory tracking performance in the Cartesian space. The inner-loop controller can not track the reference trajectory.
V-A Tracking Control
We compare the tracking performance of ours, with [2]. The reason is that we share the same minimal assumptions about the inner-loop controller, which, to the best of our knowledge, are the most general in the literature.
V-A1 Scenario 1: Normal Case without Payload
The robot tracks a reference trajectory without a payload. The tracking performance of the learning-based adaptive controller [2] and ours is shown in Fig. 5. Fig. 5a compares the reference (dashed) and actual joint trajectories of $q_{1}$, $q_{2}$, and $q_{3}$, where blue and red lines represent our method and [2], respectively. The x-axis denotes time (s), and the y-axis denotes joint angles (rad). Fig. 5b shows the Cartesian trajectories in $x$, $y$, and $z$. The close alignment of the final and reference trajectories indicates that both controllers achieve accurate tracking despite model uncertainties and an imperfect inner loop.
The proposed method achieves tracking performance comparable to the learning-based adaptive controller in the no-payload experiment. As shown in Fig. 6, both approaches yield similarly smooth control signals. Transient responses occur initially, particularly in the second joint (Fig. 5a), due to the offset between the initial and reference positions. The learning-based controller exhibits a longer transient phase (see Figs. 5a and 6). This reflects a typical trade-off in adaptive control between transient duration and tracking accuracy.
(a) Joint space comparison
(b) Task space comparison
Figure 5: Comparative tracking precision analysis: (a) Root Mean squared error (RMSE) in joint space: Proposed ($4.1\times 10^{-3}$, $2.76\times 10^{-2}$, $7.6\times 10^{-3}$) radians vs. [2] ($4.3\times 10^{-3}$, $2.7\times 10^{-2}$, $5.7\times 10^{-3}$) radians for each joint. (b) Cartesian tracking error RMSE: Proposed ($2.3$ mm, $2.4$ mm, $15.5$ mm) vs. [2] ($2.9$ mm, $2.1$ mm, $16.1$ mm) for each joint. Data sampled at 1 kHz over a 10 s trajectory execution.
Figure 6: Inner-loop control command comparison for $10$ s. $u_{1}$, $u_{2}$, and $u_{3}$ are the control inputs (voltages) for joint $q_{1}$, $q_{2}$, and $q_{3}$, respectively.
V-A2 Scenario 2: Robustness Testing with Different Payloads
We evaluate controller robustness with payloads of 1 kg, 1.5 kg, and 2 kg attached to the end-effector. Each payload experiment is repeated five times for ANOVA analysis, with identical controller and observer parameters. Fig. 7a and Fig. 7b present the tracking error ANOVA results for both methods. As shown earlier, our approach achieves a shorter transient time than the learning-based adaptive controller. For clarity, transient data (first 5 s) are excluded, and the plots show steady-state tracking errors (after 5 s).
(a) ANOVA analysis for ours
(b) ANOVA analysis for learning-based adaptive control [2]
Figure 7: Steady-state tracking error comparison under different payloads. (a) ANOVA of ours. (b) ANOVA of the learning-based adaptive control [2]. The proposed method shows consistent tracking across payloads, as indicated by uniform box plots, while the adaptive method exhibits larger error variations, implying reduced robustness. Significance levels: ns ($p>0.05$), * ($p\leq 0.05$), ** ($p\leq 0.01$), *** ($p\leq 0.001$). Each experiment is repeated five times.
As shown in Fig. 7a, the ANOVA results indicate that our method maintains statistically consistent errors across different payloads ($p>0.05$), demonstrating strong robustness, whereas the adaptive method shows significant variations with payload changes ($p<0.05$).
In summary, our approach offers several advantages over the learning-based adaptive control [2]. As shown in Fig. 5, it achieves comparable tracking performance without a payload, lower steady-state error, and stable performance under varying payloads (Fig. 7). It also exhibits a shorter transient response, as seen in the trajectories and control signals during the first second (Figs. 5 and 6). Moreover, our design is simpler than a neural network, requiring fewer tunable parameters and offering greater interpretability. Finally, the proposed disturbance compensation supports both tracking and safe control with state constraints, whereas [2] cannot handle constraints.
V-B Safe Control
It is worth noting that $\gamma$ in (53) critically influences the CBF’s behavior. A larger $\gamma$ drives the trajectory closer to the safety set boundary (more aggressive), while a smaller $\gamma$ keeps it farther away (more conservative). To achieve a safe yet efficient controller, we set $\gamma=10$ in all experiments.
We have conducted experiments in three scenarios: (1) Safety and Robustness Testing with Different Payloads, (2) Safety Testing Subject to External Disturbance, and (3) Safety Testing Subject to Gravity Disturbance. The first scenario will be described in detail, whereas the second and third scenarios are demonstrated in our accompanying video.
A virtual boundary is defined at $y_{0}=\begin{bmatrix}y_{1}&-0.1&y_{3}\end{bmatrix}^{T}$ as a spatial constraint for the end effector. It moves freely along the $x$ and $z$ axes but is constrained along the $y$ axis with a lower limit of 0.1-0.1 m. The system is tested with payloads of $1$, $1.5$, and $2$ kg attached to the end effector.
Three approaches are compared: (1) conventional CBF with a nominal model (see Fig. 8); (2) our robust CBF with a nominal model and an ESO for disturbance estimation, compared to the DOB-CBF [16] (see Fig. 9a); and (3) our robust CBF with a nominal model, ESO for disturbance estimation, and estimation error bound, compared to the DOB-CBF with estimation error bound (see Fig. 9b).
Figure 8: Safety boundary: $\bm{y_{0}}=\begin{bmatrix}\bm{y_{1}}&\bm{-0.1}&\bm{y_{3}}\end{bmatrix}^{T}$. Conventional CBF with nominal model: safety performance with variances under different payloads, with the lowest value of $h$ function at $5.9$ mm.
(a) Estimation error bounds not included
(b) Estimation error bounds included
Figure 9: Safety boundary: $\bm{y_{0}}=\begin{bmatrix}\bm{y_{1}}&\bm{-0.1}&\bm{y_{3}}\end{bmatrix}^{T}$. (a) CBFs with the nominal model and disturbance compensation by ESO (solid lines) and DOB (dashed lines) show consistent (robust) safety performance under different payloads, with the lowest $h$ value being 0.366-0.366 mm (unsafe) for our method and 0.597-0.597 mm (unsafe) for the DOB-CBF. (b) CBFs with the nominal model, disturbance compensation by ESO (solid lines) and DOB (dashed lines), and error bounds show consistent (robust) safety performance under different payloads, with the lowest $h$ value being 26.1 mm (safe and less conservative) for our method and 56.8 mm (safe and more conservative) for the DOB-CBF.
Fig. 8 shows the safety performance of the CBF using only the nominal second-order model (12), which neglects model errors, inner-loop uncertainty, and external disturbances. The safety function $h$ varies under different payloads, revealing limited robustness.
We implement the DOB-CBF [16] for comparison. The DOB-CBF estimates $b_{e}(x,d)=\frac{\partial\zeta}{\partial q}f$ in the CBF, while our method directly estimates $f$ in the original dynamics. Thus, the two observers target different quantities, and their gains are not directly comparable. For fairness, excluding error bounds, we tune both observers’ gains to achieve similar minimum $h$ magnitudes, indicating comparable safety performance. Fig. 9a compares the $h$ functions of the DOB-CBF and the proposed method, showing that:
  1. 1.
    The safety function $h$ exhibits consistent performance under different payloads, indicating robustness of both methods.
  2. 2.
    A slight safety violation occurs due to disturbance estimation error, with minimum values of 0.597-0.597 mm for DOB-CBF and 0.366-0.366 mm for the proposed method, highlighting the need to account for estimation error bounds in both.
We collect experimental data for $\hat{b}_{e}(x,d)$ and $\hat{f}$ from the DOB and ESO, respectively. Their finite differences yield the maximum rates of change, used as estimation error bounds. Fig. 9b shows the resulting $h$ values after incorporating these bounds. Compared with Fig. 9a, we observe that:
  1. 1.
    The trajectories remain safe and consistent across different payloads, demonstrating the robustness of both methods.
  2. 2.
    Incorporating the estimation error bound raises $h$, keeping trajectories strictly within the safety set (positive $h$ values).
  3. 3.
    With error bounds, our method stays closer to the safety boundary than [16], indicating lower conservatism. This may be because [16] estimates $b_{e}(x,d)=\frac{\partial\zeta}{\partial q}f$, which includes a state-dependent Jacobian, making it more sensitive to measurement noise and compounding estimation errors.
VI Conclusion and future work
This paper presents a unified framework that improves tracking performance and ensures safe kinematic control for closed-architecture robotic systems. The proposed method integrates seamlessly into existing cascade control structures as an add-on, enabling robust and safe control under uncertain dynamics, external disturbances, and unknown inner-loop controllers. Stability and safety are theoretically guaranteed, and real-time validation is conducted on the PUMA 500 industrial robot. Future work will extend the framework to visual servoing systems, addressing uncertainties in the vision pipeline.